# SwitchHead: Accelerating Transformers with Mixture-of-Experts Attention

### v2

@@ Abstract
The costly self-attention layers in modern Transformers require memory and compute quadratic in sequence length. Existing approximation methods usually underperform and fail to obtain significant speedups in practice. Here we present SwitchHead—a novel method that reduces both compute and memory requirements and achieves wall-clock speedup, while matching the language modeling performance of baseline Transformers with the same parameter budget. SwitchHead uses Mixture-of-Experts (MoE) layers for the value and output projections and requires 4 to 8 times fewer attention matrices than standard Transformers. Our novel attention can also be combined with MoE MLP layers, resulting in an efficient fully-MoE “SwitchAll” Transformer model. Our code is public. [^1] https://github.com/robertcsordas/moe_attention ${}^{\dagger}$ork done at IDSIA.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable capabilities (Radford et al., 2019; Brown et al., 2020; OpenAI, 2022; 2023) and great versatility (Bubeck et al., 2023). However, training enormous Transformers (Vaswani et al., 2017; Schmidhuber, 1992) requires a considerable amount of computing power and memory, which is not accessible to most researchers, academic institutions, and even companies. Even running them in inference mode, which is much less resource-intensive, requires significant engineering effort (Gerganov, 2023). Accelerating big Transformers remains an important open research question.
In the literature, Mixture of Experts (MoE)-based feedforward multi-layer perceptron (MLP) layers (Shazeer et al., 2017; Jacobs et al., 1991; Ivakhnenko & Lapa, 1965) have been popular methods to scale up Transformers to a large number of parameters (Lewis et al., 2021; Lepikhin et al., 2021; Fedus et al., 2022; Clark et al., 2022; Chi et al., 2022). However, in these works, the parameter efficiency of MoEs has not been studied; MoE models have been typically compared to dense baselines with the same number of FLOPs but with much less parameters. In fact, it has remained a common belief that they are not competitive against their dense counterparts with the same number of parameters. More recently, this parameter-efficiency perspective of MoEs has been studied by Csordás et al. (2023). They have proposed a novel MoE model, called $\sigma$-MoE, which not only speeds up Transformer training through reduced resource requirements, but also achieves performance comparable to or even superior to its parameter-matched dense counterparts; effectively demonstrating the promise of MoEs to scaling up neural networks.
Such MoE research often focuses on accelerating MLPs. However, attention layers (Schmidhuber, 1991; Bahdanau et al., 2015) also account for a considerable amount of compute and memory usage in Transformers, especially for long context sizes. Existing methods, such as linear attention (Schmidhuber, 1991; Katharopoulos et al., 2020; Choromanski et al., 2021; Peng et al., 2021; Schlag et al., 2021), typically fail to achieve significant speedups in practice (Dao et al., 2022) and often underperform its quadratic counterpart. As an alternative, MoE-based approaches for attention have been also proposed (Zhang et al., 2022; Peng et al., 2020), but in practice, they only achieve a modest reduction in computing and memory requirements, and typically require a lot of engineering tricks for successful training. Generally, MoE-based attention remains underexplored.
Here we present a novel MoE-based attention mechanism, SwitchHead, which reduces the number of attention matrices that need to be computed and stored. Our method is based on the $\sigma$-MoE by Csordás et al. (2023) and does not require regularization or extra tricks for stable training. We evaluate our method on a variety of language modeling datasets and two model sizes. We demonstrate that SwitchHead can achieve performance comparable to parameter-matched baselines, with a fraction of the compute and memory budget. Furthermore, we introduce the “SwitchAll” model, a fully MoE-based Transformer model, that combines a $\sigma$-MoE-based MLP layer with our SwitchHead attention, often outperforming dense baselines with the same parameter budgets. Lastly, we analyze the attention maps of our SwitchHead, and find that the maximum of attention maps taken over all heads are qualitatively similar to the dense baselines, indicating a significant reduction in redundancy without a loss of expressivity. Additionally, expert selections are often interpretable.
@@ 2 Method
Figure 1: Schematic representation of SwitchHead. It consists of few independent heads, each with multiple experts for value and output projections. Each head has a single attention matrix.
@@ 2.1 Background
The standard multi-head self-attention (MHA) layer (Vaswani et al., 2017) consists of four major steps: (1) computing key (K), query (Q), and value (V) projections, (2) computing the attention matrix, (3) using the attention matrix to project the values, and (4) mapping the projected values to the output. Let $h$, $T$, $d_{\text{model}}$, $d_{\text{head}}$ denote positive integers. Let ${\bm{x}}\in\mathbb{R}^{T\times d_{\text{model}}}$ denote an input to the MHA layer, $T$ be the sequence length, and $d_{\text{model}}$ denote the size of the hidden representations of the model. ${\bm{W}}_{\{K,V,Q\}}^{h}\in\mathbb{R}^{{d_{\text{model}}}\times{d_{\text{head}}}}$ are the projection matrices for head $h$. Then ${\bm{K}}^{h}={\bm{x}}{\bm{W}}_{K}^{h}$, ${\bm{Q}}^{h}={\bm{x}}{\bm{W}}_{Q}^{h}$, and ${\bm{V}}^{h}={\bm{x}}{\bm{W}}_{V}^{h}$ (thus ${\bm{K}}^{h},{\bm{Q}}^{h},{\bm{V}}^{h}\in\mathbb{R}^{T\times{d_{\text{head}}}}$) are the keys, queries, and values, respectively. The attention matrix for the head $h$, ${\bm{A}}^{h}\in\mathbb{R}^{T\times T}$, and the output ${\bm{y}}\in\mathbb{R}^{T\times{d_{\text{model}}}}$ are calculated as follows:
@@
${\bm{A}}^{h}$ $=\mathrm{softmax}\left(\frac{1}{\sqrt{d_{\text{model}}}}{{\bm{Q}}^{h}{\bm{K}}^{h\intercal}}\right)$ (1)
${\bm{y}}$ $={\bm{W}}_{O}({\bm{A}}^{0}{\bm{V}}^{0}|{\bm{A}}^{1}{\bm{V}}^{1}|...|{\bm{A}}^{H}{\bm{V}}^{H})$ (2)
@@
where |||| denotes concatenation in the last dimension, the $\mathrm{softmax}(\cdot)$ is also over the last dimension, and ${\bm{W}}_{O}\in\mathbb{R}^{d_{\text{model}}\times Hd_{\text{head}}}$. However, an alternative formulation reflects the role of ${\bm{W}}_{O}$ better. Let us divide ${\bm{W}}_{O}$ along the second dimension into submatrices for each head, ${\bm{W}}_{O}^{h}\in\mathbb{R}^{{d_{\text{model}}}\times d_{\text{head}}}$, such that ${\bm{W}}_{O}={\bm{W}}_{O}^{0}|{\bm{W}}_{O}^{1}|...|{\bm{W}}_{O}^{H}$. In this case, the output can be equivalently written as:
${\bm{y}}$ $=\sum_{h}{\bm{W}}_{O}^{h}{\bm{A}}^{h}{\bm{V}}^{h}$ (3)
@@
From this, it can be seen that all computations are local to the heads. Computing the attention matrix $A^{h}$ and the readout $A^{h}V^{h}$ requires compute in order of $O(Hd_{\text{head}}T^{2})$ MACs (multiplication-accumulation operation). During training, it requires the storage of $O(HT^{2})$ for the attention matrices and $O(HTd_{\text{head}})$ numbers for storing the sub-results of the projections. Given a sufficiently long sequence, computing the attention matrix and projecting the values will dominate the compute requirements due to the quadratic dependence on the sequence length $T$.
2.2 From Dense to SwitchHead
@@
Our goal is to obtain resource reductions while maintaining the fundamental properties of attention and retaining a fully expressive attention matrix. In fact, there is still room for improvement: modern LLMs use tens of heads (Brown et al., 2020; Touvron et al., 2023). Are so many of them all necessary? As we show later in Sec. 3, indeed, naively reducing the number of heads (while keeping the same number of parameters by increasing the head dimension) results in performance loss. Explaining the reason for the need for many heads is beyond the scope of this paper. Nevertheless, here are some hypotheses: (1) they provide multiple inputs for the operations that the network performs in each step, (2) they are specialized and provide inputs only for specific operations. In this case, each operation would use a different subset of heads. (3) They may also provide alternatives with different initializations, some being more successful than others, thus enabling better learning. Among these, some (2) and (3) offer an opportunity for resource savings: if not all heads are needed at the same time, it might be possible to switch between them. The simplest method of doing so is to produce a gating signal using a linear projection ${\bm{W}}_{S}\in\mathbb{R}^{d_{\text{model}}\times H}$, and use the ones with the highest activation, by replacing Eq. 3 with Eq. 6:
@@
$\mathcal{E}$ $=\operatorname*{arg\,topk}({\bm{s}},K),\mathcal{E}\subset\{1,...,H\}$ (5)
${\bm{y}}[t,c]$ $=\sum_{h\in\mathcal{E}}{\bm{s}}[t,h]({\bm{W}}_{O}^{h}{\bm{A}}^{h}{\bm{V}}^{h})[t,c]$ (6)
where ${\bm{y}}[t,c]$ denotes indexing the specific element of the matrix, specifically denoting timestep $t$ and channel $c$. Following Csordás et al. (2023), we use a non-competitive selection function. Intuitively, this corresponds to choosing a subset of attention heads for each output position. Our preliminary experiments confirmed that this method is indeed feasible for language modeling on WikiText-103. However, it is difficult to achieve acceleration and memory savings with this method. To see why, notice that the entries of the attention matrix ${\bm{A}}^{h}$ depend on pairs of inputs in different positions, but the choice is made only based on the output position. Thus, in the worst case, all possible projections have to be computed on the "source side" for the keys and values, which we would like to avoid.
An alternative approach, which we propose here, is to perform the conditional computation on the projections, independently for the source side ($K$ and $V$) and the destination side ($Q$ and output). This avoids conditional computation that involves the attention matrix itself. The obvious way to make the projections conditional is to use Mixture of Experts (MoEs). In this case, the concepts of "heads" are not well defined anymore. Therefore, we define a head to be a specific, computed attention matrix. For each head $h$, we define a list of $E$ experts. Then, the projection matrices become ${\bm{W}}_{K}^{h,e}$, ${\bm{W}}_{Q}^{h,e}$, ${\bm{W}}_{V}^{h,e}$ and ${\bm{W}}_{O}^{h,e}$, where $h$ denotes the head index and $e$ the specific expert. Then we compute the source-side expert selection as following:
@@
We compute the destination-side experts similarly: ${\bm{s}}_{D}^{h}=\sigma({\bm{x}}{\bm{W}}_{D}^{h})$, $\mathcal{E}_{D}^{h}=\operatorname*{arg\,topk}({\bm{s}}_{D}^{h},K),\mathcal{E}_{S}^{h}\subset\{1,...,E\}$. Then, the value projection ${\bm{V}}^{h}$ is computed as a weighted sum of the selected experts:
@@
As we’ll show, it is not necessary to make all projections MoEs. In Section 3.1 we show that keeping a single copy of the projections $Q$ and $K$ and reusing them for all experts is beneficial. We call this method SwitchHead. If this method can reduce the number of heads $H$ by having more experts, $E$, then it provides an easy way to reduce the resource requirements of MHA. Note that our method does not depend on the specific implementation of the attention, allowing easy experimentation and research. A schematic representation is shown in Fig. 1.
Unlike standard MoE methods, we found that no regularization is necessary to achieve good performance with our method.
2.3 Resource Usage of Different Methods
In this section, we discuss the compute and memory usage of different attention variants. We will define the compute in terms of the number of multiply-accumulate operations (MACs, also used by Zhang et al. (2022)), which is arguably better defined than FLOPs (e.g., does one step of the matrix multiplication count as 1 FLOP or 2? Do we include the softmax?). All calculations will be presented for a single attention layer for a single sequence, and they are presented this way in all our tables. Both the memory and compute requirements scale linearly with both the batch size and the number of layers.
Consider a sequence of inputs of length $T$, with representation size $d_{\text{model}}$. Let $d_{\text{head}}$ be the width of the $K$, $Q$, and $V$ projections used for the attention layer. For Transformer XL-style attention, let the size of the context be $CT$, where $C-1$ is the number of past chunks included in the context of the current attention step. We can divide the computation into two major parts: calculating the projections, which do not involve the attention map, and calculating the attention map and projecting the sequence of values using it.
First, consider the case of the standard Transformer XL (Dai et al., 2019). Here, from the input ${\bm{x}}\in\mathbb{R}^{T\times d_{\text{model}}}$, we calculate the ${\bm{K}}^{h},{\bm{Q}}^{h},{\bm{V}}^{h}\in\mathbb{R}^{T\times{d_{\text{head}}}}$ using projection matrices of shape $\mathbb{R}^{d_{\text{model}}\times d_{\text{head}}}$. The output after the attention is projected in a similar manner (Eq. 3). Thus, the projections take a total of $4Td_{\text{model}}d_{\text{head}}$ MACs per head. For backpropagation, we have to store all the intermediate results. This takes $Td_{\text{head}}$ numbers of ${\bm{K}}^{h}$, ${\bm{Q}}^{h}$ and ${\bm{V}}^{h}$. Also, the projected values should be stored. They have an identical shape, therefore, the total memory used by projections is $4Td_{\text{head}}$ numbers per head. Now consider the resource usage related to the attention matrix. It involves calculating the product of ${\bm{Q}}^{h}{\bm{K}}^{h\intercal}$, which takes ${d_{\text{head}}}CT^{2}$ MACs (multiplication by $C$ is needed because the shape of ${\bm{K}}^{h}$ and ${\bm{V}}^{h}$ for Transformer XL is $CT\times d_{\text{head}}$). The projection of the values with the attention matrix ${\bm{A}}^{h}{\bm{V}}^{h}$ is similar. For the memory usage, the attention needs $CT^{2}$ numbers, but it needs to be stored both before and after the activation function. In addition, calculating the projection of the position encodings is necessary. This depends on the implementation, but in our case, it involves a matrix multiplication, and the total amount of computation is $2{d_{\text{head}}}{d_{\text{model}}}TC$, and it needs $2{d_{\text{head}}}TC$ numbers of storage. Thus the resource requirements are:
$N_{\text{MAC}}^{\text{XL}}$ $=H\left(4Td_{\text{head}}d_{\text{model}}+2CT^{2}d_{\text{head}}+2CTd_{\text{head}}d_{\text{model}}\right)$ (11)
$N_{\text{mem}}^{\text{XL}}$ $=H\left(4Td_{\text{head}}+2CT^{2}+2CTd_{\text{head}}\right)$ (12)
The resource usage of SwitchHead is different. First, the number of heads $H$ is significantly reduced, but $d_{\text{head}}$ is typically larger. Additionally, there are $K$ experts active at the same time. Here, we only consider the case where the value and outputs are experts, but ${\bm{Q}}^{h}$ and ${\bm{K}}^{h}$ are not (this version performs the best; see Sec. 3.1). Then, we have two projections that are identical with that of Transformer XL, and two MoE-based projections. These use $TKd_{\text{model}}d_{\text{head}}$ MACs to calculate the projection and another $TKd_{\text{head}}$ to calculate their weighted average. With a smart kernel implementation, memory usage is not affected by $K$, thus the formula remains the same as Eq. 12 (note, however, that $H$ and $d_{\text{head}}$ are very different in practice). The compute requirement can be calculated as:
$N_{\text{MAC}}^{\text{{SwitchHead}}}$ $=H\left(2Td_{\text{head}}d_{\text{model}}+2TKd_{\text{head}}(d_{\text{model}}+1)+2CT^{2}d_{\text{head}}+2CTd_{\text{head}}d_{\text{model}}\right)$ (13)
Additionally, the expert selection logic needs minimal additional resources, which can be ignored. Note that the comparison between the MACs of the standard (Eq. 11) and SwitchHead (Eq. 13) depends on the exact values of the hyper-parameters. However, as we’ll see in Sec. 3, in our typical configurations, SwitchHead provides good predictive performance with significantly lower $H$ compared to the standard Transformer, resulting in reduced resource usage in the end.

### v3

@@ Abstract
22footnotetext: Work done at IDSIA.
Despite many recent works on Mixture of Experts (MoEs) for resource-efficient Transformer language models, existing methods mostly focus on MoEs for feedforward layers. Previous attempts at extending MoE to the self-attention layer fail to match the performance of the parameter-matched baseline. Our novel SwitchHead is an effective MoE method for the attention layer that successfully reduces both the compute and memory requirements, achieving wall-clock speedup, while matching the language modeling performance of the baseline Transformer. Our novel MoE mechanism allows SwitchHead to compute up to 8 times fewer attention matrices than the standard Transformer. SwitchHead can also be combined with MoE feedforward layers, resulting in fully-MoE “SwitchAll” Transformers. For our 262M parameter model trained on C4, SwitchHead matches the perplexity of standard models with only 44% compute and 27% memory usage. Zero-shot experiments on downstream tasks confirm the performance of SwitchHead, e.g., achieving more than 3.5% absolute improvements on BliMP compared to the baseline with an equal compute resource. [^1] Our code is public: https://github.com/robertcsordas/switchhead
@@ 1 Introduction
Figure 1: A schematic representation of SwitchHead. It consists of a few independent heads, each with multiple experts for value and output projections. Each head has a single attention matrix.
@@
Large language models (LLMs) have shown remarkable capabilities [1, 2, 3, 4] and great versatility [5]. However, training large Transformers [6, 7] requires a considerable amount of computing power and memory, which is not accessible to most researchers, academic institutions, and even companies. Even running them in inference mode—typically much less resource-intensive—requires significant engineering effort [8]. Accelerating Transformers remains an important research question.
In this context, Mixture of Experts (MoE) layers [9, 10, 11] have become popular to efficiently scale up Transformers to a large number of parameters [12, 13, 14, 15, 16, 17]. However, most of these works mainly focus on applying MoE to the 2-layer feedforward blocks [6], i.e., the multi-layer perceptron (MLP) components of the Transformer, while keeping the self-attention layers unchanged. Given that attention also accounts for a considerable amount of compute and memory usage in Transformers (especially for long context sizes), using MoE for attention has potential to further improve resource efficiency in Transformers. While MoE-based attention remains underexplored in general, there are existing works on MoE approaches for attention [18, 19]. However, in practice, previously proposed methods typically require a lot of engineering tricks for successful training, and most importantly, only achieve a modest reduction in computing and memory requirements in the end (as we also confirm in our experiments).
Here, we present a novel MoE-based attention method, SwitchHead, whose mechanism allows to reduce the number of attention matrices that need to be computed and stored. Following $\sigma$-MoE [17], our method uses a non-competitive selection activation function (sigmoid), and does not require regularization or extra tricks for stable training. Importantly, we show that it is possible to compute the MoE projections outside of the attention core, which enables a significant reduction in the number of computed attention maps, resulting in significant resource savings. Our thorough investigation shows that it is enough to choose the value and output projections from a pool of experts and share keys and queries between them.
We evaluate our method on C4 [20], Enwik8 [21], peS2o [22] and Wikitext 103 [23], with two model sizes (47M and 262M). Additionally, we measure the zero-shot performance of our main models on Lambada [24], BLiMP [25], and Children’s Books Test [26] datasets. Our experiments demonstrate that SwitchHead can achieve performance comparable to parameter-matched baselines with just a fraction of the compute and memory budget. In addition, we introduce “SwitchAll”, a fully MoE-based Transformer model, that combines a $\sigma$-MoE-based MLP layer with our SwitchHead attention, often outperforming dense baselines with the same parameter budgets.
Finally, we analyze the attention maps of our SwitchHead. We find that the attention maps taken over all heads are qualitatively similar to the dense baselines, indicating a significant reduction in redundancy without a loss of expressivity. In addition, expert selections are often interpretable.
@@ 2.1 Background
The standard multi-head self-attention (MHA) layer [6] consists of four major steps: (1) compute key, query, and value projections, (2) compute the attention matrix, (3) use the attention matrix to project the values, and (4) map the projected values to the output. Let $h$, $T$, $n_{\text{heads}}$, $d_{\text{model}}$, $d_{\text{head}}$ denote positive integers. Let ${\bm{x}}\in\mathbb{R}^{T\times d_{\text{model}}}$ denote an input to the MHA layer with $n_{\text{heads}}$ heads, $T$ be the sequence length, and $d_{\text{model}}$ denote the size of the hidden representations of the model. ${\bm{W}}_{\{K,V,Q\}}^{h}\in\mathbb{R}^{{d_{\text{model}}}\times{d_{\text{head}}}}$ are the projection matrices for head $h\in\{1,...,n_{\text{heads}}\}$. Then ${\bm{K}}^{h}={\bm{x}}{\bm{W}}_{K}^{h}$, ${\bm{Q}}^{h}={\bm{x}}{\bm{W}}_{Q}^{h}$, and ${\bm{V}}^{h}={\bm{x}}{\bm{W}}_{V}^{h}$ (thus ${\bm{K}}^{h},{\bm{Q}}^{h},{\bm{V}}^{h}\in\mathbb{R}^{T\times{d_{\text{head}}}}$) are the keys, queries, and values, respectively. The attention matrix for the head $h$, ${\bm{A}}^{h}\in\mathbb{R}^{T\times T}$, and the output ${\bm{y}}\in\mathbb{R}^{T\times{d_{\text{model}}}}$ are calculated as follows:
@@
${\bm{A}}^{h}$ $=\mathrm{softmax}\left(\frac{1}{\sqrt{d_{\text{head}}}}{{\bm{Q}}^{h}{{\bm{K}}^{h}}^{\intercal}}\right)$ (1)
${\bm{y}}$ $=({\bm{A}}^{1}{\bm{V}}^{1}|{\bm{A}}^{2}{\bm{V}}^{2}|...|{\bm{A}}^{n_{\text{heads}}}{\bm{V}}^{n_{\text{heads}}}){\bm{W}}_{O}$ (2)
@@
where |||| denotes concatenation in the last dimension, the $\mathrm{softmax}(\cdot)$ is also over the last dimension, and ${\bm{W}}_{O}\in\mathbb{R}^{n_{\text{heads}}d_{\text{head}}\times d_{\text{model}}}$. However, an alternative formulation reflects the role of ${\bm{W}}_{O}$ better. Let us divide ${\bm{W}}_{O}$ along the first dimension into submatrices for each head, ${\bm{W}}_{O}^{h}\in\mathbb{R}^{{d_{\text{head}}}\times{d_{\text{model}}}}$, such that ${\bm{W}}_{O}=\left({{\bm{W}}_{O}^{1}}^{\intercal}|{{\bm{W}}_{O}^{2}}^{\intercal}|...|{{\bm{W}}_{O}^{n_{\text{heads}}}}^{\intercal}\right)^{\intercal}$. In this case, the output (Eq. 2) can be equivalently written as:
${\bm{y}}$ $=\sum_{h}{\bm{A}}^{h}{\bm{V}}^{h}{\bm{W}}_{O}^{h}$ (3)
@@
From this, it can be seen that all computations are local to each head. Computing the attention matrix ${\bm{A}}^{h}$ and the readout ${\bm{A}}^{h}{\bm{V}}^{h}$ requires compute in order of $O(n_{\text{heads}}d_{\text{head}}T^{2})$ MACs (multiplication-accumulation operation [^2] The number of MACs is a metric used in prior work [18], which is independent of both the specific hardware and implementation, unlike wall-clock time. For wall-clock-time measurements, see Sec. 3.7.). During training, it requires the storage of $O(n_{\text{heads}}T^{2})$ for the attention matrices and $O(n_{\text{heads}}Td_{\text{head}})$ for storing the sub-results of the projections. Given a sufficiently long sequence, computing the attention matrix and projecting the values will dominate the compute requirements due to the quadratic dependence on the sequence length $T$.
2.2 From Dense to SwitchHead Attention Layer
@@
Our goal is to obtain resource reductions while maintaining the fundamental properties of attention and retaining a fully expressive attention matrix. For that, we start from the following observation: modern LLMs use tens of heads [2, 27]. Are so many of them all necessary? As we show later in Sec. 3, indeed, naively reducing the number of heads (while keeping the same number of parameters by increasing the head dimension) results in performance loss. Explaining the reason for the need for many heads is beyond the scope of this paper. Nevertheless, here are some hypotheses: (1) they provide multiple inputs for the operations that the network performs in each step, (2) they are specialized and provide inputs only for specific operations (in this case, each operation would use a different subset of heads), (3) they may provide diverse outputs due to different initializations, some being more successful than others, thus enabling better learning. Among these, (2) and (3) may offer an opportunity for resource savings: if not all heads are needed at the same time, it might be possible to switch among them depending on the context.
One naive method to achieve this is to use a gating signal using a linear projection ${\bm{W}}_{S}\in\mathbb{R}^{d_{\text{model}}\times n_{\text{heads}}}$, and use the heads with the highest score, by replacing Eq. 3 with Eq. 6:
@@
$\mathcal{E}$ $=\operatorname*{arg\,topk}({\bm{s}},k),\mathcal{E}\subset\{1,...,n_{\text{heads}}\}$ (5)
${\bm{y}}[t,c]$ $=\sum_{h\in\mathcal{E}}{\bm{s}}[t,h]({\bm{A}}^{h}{\bm{V}}^{h}{\bm{W}}_{O}^{h})[t,c]$ (6)
where ${\bm{y}}[t,c]\in\mathbb{R}$ denotes indexing the specific element of the output matrix ${\bm{y}}\in\mathbb{R}^{T\times{d_{\text{model}}}}$, for timestep $t$ and channel $c$, and $k$ is the number of active experts. Following the $\sigma$-MoE method [17], we use a non-competitive selection function (sigmoid $\sigma$ in Eq. 4). Now, let us define the source side of attention as the keys and values and the destination side as the queries and output. Intuitively, the above method corresponds to choosing a subset of attention heads based on the destination side alone [^3] To clarify, we allocate a routing function for each of key/value/query projections; these routing functions belong to the source or destination side accordingly. If we compare Eq. 10 and Eq. 6, one can notice that the routing function in Eq. 6 effectively corresponds to what we define as the destination-side routing in Eq. 10.. Our preliminary experiments confirmed that this method is indeed feasible for language modeling on WikiText-103. However, it is difficult to achieve acceleration and memory savings with this method. To see why, notice that the entries of the attention matrix ${\bm{A}}^{h}$ depend on pairs of tokens, one for the source and one for the destination side, but the choice is made only based on the destination side. Thus, in the worst case, for each destination, a different source might be chosen, in which case all possible source projections have to be computed for the keys and values, which we would like to avoid.
Alternatively, we propose to improve the method above by introducing conditional computations for the source and destination projections independently of each other. That is, we parameterize each of key, query, value, output projection by an independent MoE. This avoids conditional computations that involve the attention matrix itself. Our solution implements this using Mixtures of Experts (MoEs). The concepts of "heads" are no longer well defined in the conventional sense: we redefine a head as an instance of a computed attention matrix. We call the total number of them $n_{\text{heads}}$. For each head $h$, we define a separate list of $E$ experts. The total number of experts is then $n_{\text{heads}}\cdot E$. Then, the projection matrices become ${\bm{W}}_{K}^{h,e}$, ${\bm{W}}_{Q}^{h,e}$, ${\bm{W}}_{V}^{h,e}$ and ${\bm{W}}_{O}^{h,e}\in\mathbb{R}^{d_{\text{head}}\times d_{\text{model}}}$, where $h$ denotes the head index and $e$ the specific expert. Then we compute the source-side expert selection as follows:
@@
where ${\bm{W}}_{S}^{h}\in\mathbb{R}^{d_{\text{model}}\times E}$. We compute the destination-side experts similarly: ${\bm{s}}_{D}^{h}=\sigma({\bm{x}}{\bm{W}}_{D}^{h})$, $\mathcal{E}_{D}^{h}=\operatorname*{arg\,topk}({\bm{s}}_{D}^{h},k),\mathcal{E}_{S}^{h}\subset\{1,...,E\},{\bm{W}}_{D}^{h}\in\mathbb{R}^{d_{\text{model}}\times E}$. Then, the value projection ${\bm{V}}^{h}$ is computed as a weighted sum of the selected experts:
@@
As we’ll show, it is not necessary to make all projections MoEs. In Section 3.1 we show that keeping a single, head-specific copy of the query and key projections and reusing them for all experts is beneficial. We call this method SwitchHead.
Essentially, SwitchHead reduces the number of attention matrices that have to be computed ($n_{\text{heads}}$) significantly, by using multiple experts per head. Note that our method does not depend on the specific implementation of the attention, allowing for easy experimentation and research. A schematic representation is shown in Figure 1.
Table 1: Performance of SwitchHead compared to different MoA variants. MoA can outperform the baseline, but only at a price of using significantly more compute and memory. Also, SwitchHead outperforms the baseline dense Transformer. These results are on Wikitext 103. Table sorted by model perplexity.
#total params Model $n_{\text{heads}}$ Perplexity $\downarrow$ MACs Mem (floats)
47M SwitchHead 2 12.27 170.4M 0.8M
Transformer 10 12.31 453.4M 3.5M
MoA 4 12.60 223.5M 1.3M
MoA 6 12.64 306.8M 1.9M
MoA 8 12.77 390.2M 2.6M
MoA 2 12.84 140.1M 0.7M
262M MoA 8 9.50 2.9G 9.9M
SwitchHead 2 9.55 2.0G 2.9M
Transformer 16 9.66 5.4G 21.0M
MoA 12 9.68 4.1G 14.7M
MoA 4 9.69 1.7G 5.1M
MoA 2 9.87 1.1G 2.7M